OpenStudy (babyslapmafro):

Please help me determine whether the given function is homogeneous. f(x,y)=cos(x^2/(x+y))

OpenStudy (babyslapmafro):

\[f(x,y)=\cos(\frac{ x^2 }{ x+y })\]

OpenStudy (babyslapmafro):

\[f(tx,ty)=\cos(\frac{ t^2x^2 }{ tx+ty })\]

OpenStudy (babyslapmafro):

\[=\cos(t \frac{ x^2 }{ x+y })\]

OpenStudy (babyslapmafro):

is it 'not homogeneous' because you can't factor the (t) out of cosine?

OpenStudy (babyslapmafro):

f(tx,ty) does not equal (t^n)f(x,y) making it not homogeneous... correct?